# Communication Efficient LLM Pre-training with SparseLoCo

### v1

@@ Appendix D Hyperparameter Selection Details
| DiLoCo - Nesterov Outer | H=15 | |
@@
Appendix E Compression of indices in
In methods, the indices of the selected values need to be transmitted alongside the values. When values are aggressively quantized (as in SparseLoCo), this index-transmission overhead becomes significant. In SparseLoCo, we utilize chunk sizes of , so, naively, we can transmit indices in bits per transmitted value. However, with 2-bit quantization, this overhead becomes significant, motivating further index compression. Assuming a chunk size of and selection, we observe that the information-theoretic limit is bits. For practical cases considered in this work ( and ), this corresponds to 8.3, 6.3, and 5.3 bits per transmitted value, respectively. In practice, we were able to design a custom compression algorithm based on sub-chunking and coding that achieves 8.9, 6.6, and 5.6 bits per value for these cases.

### v2

@@ Appendix H Hyperparameter Selection Details
| DiLoCo - Nesterov Outer | H=15, | |
@@
Appendix I Equivalence of LOM and Global Outer Momentum
We show that DiLoCo-LOM iterates are actually equivalent to DiLoCo.
Proposition 1.
Suppose identical initialization for all , and fixed outer-momentum coefficient . Then, for all ,
where denotes the average of local momentum buffers, the averaged LOM Nesterov direction, and and the global momentum and Nesterov direction in DiLoCo, respectively. Consequently, the parameter updates of DiLoCo-LOM and DiLoCo are identical at every time step.
Proof.
We first show by induction, then obtain by linearity.
Base case (). With and , we have .
Inductive step. Assume for some . Averaging the local recursion,
By the global recursion, , hence . For the Nesterov directions,